# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v2

@@ Appendix A The Leaderboards
Rank Model Score
  1 XComposer-VL 80.13
2 InstructBlip 72.36
3 Blip2 70.03
4 Shikra-VQA 64.03
5 Shikra 62.97
6 Otter 62.22
7 Qwen-VL-Chat 61.13
8 LLAVA-1.5-13B 60.17
9 LLAVA-1.5-7B 53.81
10 MiniGPT-4 42.52
(a) Clean-Movie
Rank Model Score
  1 XComposer-VL 95.89
2 InstructBlip 95.44
3 Blip2 94.64
4 Shikra-VQA 88.25
5 Qwen-VL-Chat 79.75
6 Shikra 77.65
7 Otter 69.10
8 LLAVA-1.5-13B 68.82
9 LLAVA-1.5-7B 62.77
10 MiniGPT-4 46.98
(b) Clean-Action
Rank Model Score
  1 XComposer-VL 75.55
2 InstructBlip 63.42
3 Otter 63.04
4 Shikra-VQA 60.10
5 Blip2 59.14
6 Shikra 58.55
7 Qwen-VL-Chat 58.27
8 LLAVA-1.5-13B 56.69
9 LLAVA-1.5-7B 53.48
10 MiniGPT-4 40.23
(c) Clean-TV Show
Rank Model Score
  1 XComposer-VL 79.92
2 InstructBlip 77.84
3 Blip2 76.88
4 Shikra 73.55
5 Shikra-VQA 73.07
6 Qwen-VL-Chat 65.12
7 LLAVA-1.5-13B 62.86
8 Otter 60.95
9 LLAVA-1.5-7B 57.74
10 MiniGPT-4 42.06
(d) Clean-Profession
Rank Model Score
  1 InstructBlip 96.90
2 Blip2 96.64
3 XComposer-VL 96.43
4 Shikra-VQA 82.85
5 Shikra 80.17
6 Qwen-VL-Chat 75.84
7 LLAVA-1.5-13B 75.12
8 LLAVA-1.5-7B 69.74
9 Otter 62.50
10 MiniGPT-4 52.81
(e) Clean-Landmark
Rank Model Score
  1 XComposer-VL 75.00
2 Qwen-VL-Chat 63.11
3 InstructBlip 62.66
4 Otter 59.56
5 Blip2 59.50
6 LLAVA-1.5-13B 53.88
7 Shikra-VQA 53.31
8 Shikra 52.84
9 LLAVA-1.5-7B 47.80
10 MiniGPT-4 38.98
(f) Clean-Anime
Rank Model Score
  1 XComposer-VL 85.96
2 InstructBlip 83.80
3 Blip2 78.09
4 Shikra-VQA 69.83
5 Qwen-VL-Chat 69.58
6 Shikra 67.43
7 LLAVA-1.5-13B 61.20
8 Otter 58.33
9 LLAVA-1.5-7B 48.80
10 MiniGPT-4 40.86
(g) Clean-Clothes
Rank Model Score
  1 XComposer-VL 87.91
2 InstructBlip 79.91
3 Blip2 79.23
4 Qwen-VL-Chat 68.71
5 Shikra-VQA 62.93
6 Shikra 61.81
7 LLAVA-1.5-13B 59.04
8 LLAVA-1.5-7B 56.04
9 Otter 52.97
10 MiniGPT-4 39.13
(h) Clean-Celebrity
Rank Model Score
  1 XComposer-VL 91.27
2 InstructBlip 90.28
3 Blip2 89.78
4 Shikra-VQA 80.23
5 Shikra 75.84
6 Qwen-VL-Chat 71.37
7 LLAVA-1.5-13B 68.16
8 Otter 63.22
9 LLAVA-1.5-7B 52.87
10 MiniGPT-4 47.72
(i) Clean-Food
Rank Model Score
  1 XComposer-VL 93.19
2 InstructBlip 92.62
3 Blip2 91.84
4 Shikra-VQA 80.05
5 Shikra 77.57
6 Qwen-VL-Chat 76.11
7 Otter 73.96
8 LLAVA-1.5-13B 67.61
9 LLAVA-1.5-7B 48.84
10 MiniGPT-4 47.95
(j) Clean-Plant
Rank Model Score
  1 XComposer-VL 73.12
2 InstructBlip 62.34
3 Shikra-VQA 61.42
4 Blip2 61.13
5 Shikra 61.04
6 LLAVA-1.5-13B 54.57
7 Qwen-VL-Chat 51.28
8 LLAVA-1.5-7B 46.75
9 Otter 44.58
10 MiniGPT-4 41.73
(k) Clean-Age
Rank Model Score
  1 XComposer-VL 98.83
2 Blip2 97.12
3 InstructBlip 93.86
4 LLAVA-1.5-13B 91.12
5 Shikra-VQA 90.40
6 Shikra 85.50
7 Qwen-VL-Chat 78.13
8 Otter 78.09
9 MiniGPT-4 52.86
10 LLAVA-1.5-7B 51.96
(l) Clean-Gender
Rank Model Score
  1 XComposer-VL 86.08
2 InstructBlip 85.59
3 Shikra-VQA 83.53
4 Blip2 82.59
5 Shikra 80.51
6 Qwen-VL-Chat 74.76
7 Otter 67.11
8 LLAVA-1.5-13B 64.81
9 LLAVA-1.5-7B 56.81
10 MiniGPT-4 46.46
(m) Clean-Expression
Rank Model Score
  1 XComposer-VL 78.56
2 InstructBlip 76.31
3 Blip2 73.16
4 LLAVA-1.5-13B 66.50
5 Qwen-VL-Chat 65.19
6 Shikra-VQA 64.56
7 Shikra 64.53
8 Otter 50.44
9 LLAVA-1.5-7B 47.27
10 MiniGPT-4 38.98
(n) Clean-Race
Rank Model Score
  1 XComposer-VL 97.27
2 InstructBlip 96.09
3 Blip2 95.95
4 Shikra-VQA 82.72
5 Otter 82.47
6 Qwen-VL-Chat 79.98
7 Shikra 79.44
8 LLAVA-1.5-13B 72.11
9 LLAVA-1.5-7B 50.08
10 MiniGPT-4 47.77
(o) Clean-Animal
Rank Model Score
  1 XComposer-VL 90.97
2 InstructBlip 90.00
3 Blip2 88.53
4 Shikra-VQA 82.72
5 Qwen-VL-Chat 75.73
6 Shikra 70.06
7 LLAVA-1.5-13B 67.89
8 Otter 64.91
9 LLAVA-1.5-7B 54.31
10 MiniGPT-4 51.65
(p) Clean-Object
Rank Model Score
  1 Blip2 68.38
2 InstructBlip 67.50
3 XComposer-VL 66.06
4 Shikra-VQA 64.86
5 Otter 64.37
6 Qwen-VL-Chat 63.28
7 Shikra 63.02
8 LLAVA-1.5-13B 58.84
9 LLAVA-1.5-7B 44.96
10 MiniGPT-4 43.70
(q) Clean-OCR
Rank Model Score
  1 XComposer-VL 73.94
2 Blip2 66.98
3 InstructBlip 66.77
4 Otter 63.82
5 Qwen-VL-Chat 61.80
6 Shikra 57.21
7 Shikra-VQA 56.89
8 LLAVA-1.5-13B 55.68
9 LLAVA-1.5-7B 49.05
10 MiniGPT-4 41.47
(r) Clean-Style
Rank Model Score
  1 XComposer-VL 73.27
2 Shikra-VQA 68.47
3 InstructBlip 67.84
4 Shikra 66.95
5 Blip2 66.17
6 Otter 63.23
7 Qwen-VL-Chat 60.51
8 LLAVA-1.5-13B 59.00
9 LLAVA-1.5-7B 47.64
10 MiniGPT-4 41.02
(s) Clean-Background
Rank Model Score
  1 XComposer-VL 88.50
2 InstructBlip 88.24
3 Blip2 86.78
4 Shikra-VQA 74.14
5 Shikra 73.37
6 Qwen-VL-Chat 69.72
7 LLAVA-1.5-13B 58.36
8 Otter 53.12
9 LLAVA-1.5-7B 46.91
10 MiniGPT-4 42.90
(t) Clean-Color
Table 6: Leaderboards of all Perceptual tasks under clean scenario.
Rank Model Score
  1 XComposer-VL 78.20
2 InstructBlip 71.30
3 Blip2 68.69
4 Shikra-VQA 64.19
5 Shikra 63.47
6 Otter 61.39
7 Qwen-VL 59.27
8 LLAVA-1.5-13B 58.67
9 LLAVA-1.5-7B 52.36
10 MiniGPT-4 43.02
(a) Corru.-Movie
Rank Model Score
  1 XComposer-VL 95.86
2 InstructBlip 95.09
3 Blip2 93.87
4 Shikra-VQA 88.72
5 Shikra 77.77
6 Qwen-VL 75.09
7 Otter 70.09
8 LLAVA-1.5-13B 68.34
9 LLAVA-1.5-7B 63.00
10 MiniGPT-4 47.22
(b) Corru.-Action
Rank Model Score
  1 XComposer-VL 72.86
2 InstructBlip 62.42
3 Otter 60.71
4 Shikra 60.09
5 Shikra-VQA 59.58
6 Blip2 58.16
7 Qwen-VL 57.95
8 LLAVA-1.5-13B 55.18
9 LLAVA-1.5-7B 52.34
10 MiniGPT-4 38.70
(c) Corru.-TV Show
Rank Model Score
  1 XComposer-VL 78.48
2 InstructBlip 77.84
3 Blip2 75.77
4 Shikra 73.72
5 Shikra-VQA 73.38
6 Qwen-VL 64.97
7 Otter 62.71
8 LLAVA-1.5-13B 62.23
9 LLAVA-1.5-7B 56.95
10 MiniGPT-4 43.22
(d) Corru.-Profession
Rank Model Score
  1 InstructBlip 96.33
2 Blip2 96.28
3 XComposer-VL 95.97
4 Shikra-VQA 83.01
5 Shikra 82.38
6 LLAVA-1.5-13B 74.63
7 Qwen-VL 73.55
8 LLAVA-1.5-7B 68.78
9 Otter 61.65
10 MiniGPT-4 49.55
(e) Corru.-Landmark
Rank Model Score
  1 XComposer-VL 72.28
2 InstructBlip 62.79
3 Qwen-VL 60.44
4 Blip2 58.67
5 Otter 57.24
6 LLAVA-1.5-13B 53.67
7 Shikra-VQA 52.28
8 Shikra 52.00
9 LLAVA-1.5-7B 47.87
10 MiniGPT-4 40.91
(f) Corru.-Anime
Rank Model Score
  1 XComposer-VL 84.87
2 InstructBlip 82.84
3 Blip2 76.69
4 Shikra-VQA 70.59
5 Shikra 67.05
6 Qwen-VL 64.26
7 LLAVA-1.5-13B 60.88
8 Otter 59.87
9 LLAVA-1.5-7B 48.12
10 MiniGPT-4 41.38
(g) Corru.-Clothes
Rank Model Score
  1 XComposer-VL 87.25
2 InstructBlip 79.23
3 Blip2 78.36
4 Qwen-VL 64.89
5 Shikra-VQA 62.55
6 Shikra 61.68
7 LLAVA-1.5-13B 58.64
8 LLAVA-1.5-7B 56.36
9 Otter 53.59
10 MiniGPT-4 38.80
(h) Corru.-Celebrity
Rank Model Score
  1 XComposer-VL 90.61
2 InstructBlip 90.08
3 Blip2 89.19
4 Shikra-VQA 80.55
5 Shikra 75.63
6 Qwen-VL 70.71
7 LLAVA-1.5-13B 67.46
8 Otter 65.06
9 LLAVA-1.5-7B 52.94
10 MiniGPT-4 47.76
(i) Corru.-Food
Rank Model Score
  1 XComposer-VL 92.45
2 InstructBlip 92.23
3 Blip2 91.15
4 Shikra-VQA 80.80
5 Shikra 77.69
6 Otter 73.42
7 Qwen-VL 72.01
8 LLAVA-1.5-13B 67.22
9 LLAVA-1.5-7B 48.84
10 MiniGPT-4 48.04
(j) Corru.-Plant
Rank Model Score
  1 XComposer-VL 73.04
2 InstructBlip 61.59
3 Shikra-VQA 61.09
4 Shikra 60.40
5 Blip2 60.03
6 Qwen-VL 55.26
7 LLAVA-1.5-13B 53.67
8 LLAVA-1.5-7B 46.76
9 Otter 45.09
10 MiniGPT-4 40.84
(k) Corru.-Age
Rank Model Score
  1 XComposer-VL 98.53
2 Blip2 97.76
3 InstructBlip 93.75
4 LLAVA-1.5-13B 91.09
5 Shikra-VQA 90.17
6 Shikra 85.91
7 Otter 80.22
8 Qwen-VL 67.00
9 MiniGPT-4 52.75
10 LLAVA-1.5-7B 52.27
(l) Corru.-Gender
Rank Model Score
  1 XComposer-VL 86.24
2 InstructBlip 84.95
3 Shikra-VQA 83.56
4 Blip2 81.56
5 Shikra 80.56
6 Qwen-VL 70.55
7 Otter 66.64
8 LLAVA-1.5-13B 64.94
9 LLAVA-1.5-7B 56.16
10 MiniGPT-4 46.55
(m) Corru.-Expression
Rank Model Score
  1 XComposer-VL 78.80
2 InstructBlip 76.19
3 Blip2 71.72
4 LLAVA-1.5-13B 65.33
5 Shikra-VQA 64.73
6 Shikra 63.94
7 Qwen-VL 63.55
8 Otter 49.73
9 LLAVA-1.5-7B 47.92
10 MiniGPT-4 39.65
(n) Corru.-Race
Rank Model Score
  1 XComposer-VL 97.06
2 InstructBlip 95.86
3 Blip2 95.69
4 Shikra-VQA 83.03
5 Otter 82.84
6 Shikra 79.44
7 LLAVA-1.5-13B 71.47
8 Qwen-VL 71.43
9 LLAVA-1.5-7B 50.67
10 MiniGPT-4 48.66
(o) Corru.-Animal
Rank Model Score
  1 XComposer-VL 90.81
2 InstructBlip 89.76
3 Blip2 88.05
4 Shikra-VQA 82.76
5 Shikra 71.49
6 Qwen-VL 67.78
7 LLAVA-1.5-13B 67.23
8 Otter 66.18
9 LLAVA-1.5-7B 54.09
10 MiniGPT-4 51.02
(p) Corru.-Object
Rank Model Score
  1 Blip2 94.06
2 InstructBlip 90.54
3 XComposer-VL 87.19
4 Shikra-VQA 80.80
5 Otter 78.70
6 Shikra 77.17
7 Qwen-VL 69.66
8 LLAVA-1.5-13B 68.58
9 LLAVA-1.5-7B 49.10
10 MiniGPT-4 44.63
(q) Corru.-OCR
Rank Model Score
  1 XComposer-VL 72.34
2 InstructBlip 65.06
3 Blip2 64.86
4 Otter 61.14
5 Qwen-VL 60.34
6 Shikra 56.61
7 Shikra-VQA 55.81
8 LLAVA-1.5-13B 53.65
9 LLAVA-1.5-7B 47.67
10 MiniGPT-4 40.81
(r) Courruption-Style
Rank Model Score
  1 XComposer-VL 72.53
2 Shikra-VQA 68.25
3 InstructBlip 67.48
4 Shikra 67.19
5 Blip2 64.69
6 Otter 62.99
7 Qwen-VL 60.50
8 LLAVA-1.5-13B 58.50
9 LLAVA-1.5-7B 47.12
10 MiniGPT-4 40.53
(s) Corru.-Background
Rank Model Score
  1 XComposer-VL 88.24
2 InstructBlip 88.09
3 Blip2 86.25
4 Shikra-VQA 74.14
5 Shikra 73.07
6 Qwen-VL 70.78
7 LLAVA-1.5-13B 57.70
8 Otter 56.14
9 LLAVA-1.5-7B 46.76
10 MiniGPT-4 42.86
(t) Corru.-Color
Table 7: Leaderboards of all Perceptual tasks under corruption scenario.
Rank Model Score
  1 XComposer-VL 67.09
2 Blip2 52.55
3 LLAVA-1.5-13B 46.74
4 LLAVA-1.5-7B 45.24
5 Shikra 44.84
6 Shikra-VQA 44.36
7 Qwen-VL 44.30
8 InstructBlip 43.06
9 MiniGPT-4 40.26
10 Otter 37.97
(a) Pr.Att.-Movie
Rank Model Score
  1 XComposer-VL 86.24
2 Blip2 79.13
3 InstructBlip 75.94
4 Shikra-VQA 71.36
5 Qwen-VL 65.26
6 Shikra 58.82
7 LLAVA-1.5-13B 57.66
8 LLAVA-1.5-7B 51.62
9 MiniGPT-4 47.07
10 Otter 39.77
(b) Pr.Att.-Action
Rank Model Score
  1 XComposer-VL 55.08
2 LLAVA-1.5-7B 43.55
3 Qwen-VL 40.72
4 LLAVA-1.5-13B 40.09
5 Blip2 39.71
6 Shikra 39.01
7 MiniGPT-4 37.40
8 Shikra-VQA 36.50
9 InstructBlip 32.17
10 Otter 31.48
(c) Pr.Att.-TV Show
Rank Model Score
  1 XComposer-VL 67.14
2 Blip2 59.62
3 Shikra 55.70
4 InstructBlip 54.52
5 Shikra-VQA 54.25
6 Qwen-VL 48.66
7 LLAVA-1.5-13B 47.49
8 LLAVA-1.5-7B 42.86
9 MiniGPT-4 42.22
10 Otter 38.12
(d) Pr.Att.-Profession
Rank Model Score
  1 XComposer-VL 89.81
2 Blip2 84.85
3 InstructBlip 76.78
4 Shikra-VQA 68.28
5 LLAVA-1.5-13B 65.98
6 Shikra 64.47
7 Qwen-VL 64.29
8 LLAVA-1.5-7B 56.39
9 MiniGPT-4 49.58
10 Otter 43.89
(e) Pr.Att.-Landmark
Rank Model Score
  1 XComposer-VL 59.72
2 Blip2 45.05
3 Qwen-VL 44.95
4 LLAVA-1.5-7B 42.23
5 LLAVA-1.5-13B 40.95
6 InstructBlip 38.26
7 Shikra-VQA 38.25
8 MiniGPT-4 37.27
9 Shikra 37.02
10 Otter 36.16
(f) Pr.Att.-Anime
Rank Model Score
  1 XComposer-VL 76.43
2 Blip2 66.53
3 InstructBlip 62.14
4 Qwen-VL 51.56
5 LLAVA-1.5-13B 49.63
6 Shikra-VQA 49.02
7 Shikra 47.23
8 LLAVA-1.5-7B 42.22
9 MiniGPT-4 39.66
10 Otter 34.67
(g) Pr.Att.-Clothes
Rank Model Score
  1 XComposer-VL 73.48
2 Blip2 57.30
3 Qwen-VL 52.34
4 LLAVA-1.5-7B 45.11
5 InstructBlip 43.36
6 Shikra 43.13
7 Shikra-VQA 42.01
8 LLAVA-1.5-13B 41.59
9 MiniGPT-4 38.64
10 Otter 33.95
(h) Pr.Att.-Celebrity
Rank Model Score
  1 XComposer-VL 83.60
2 Blip2 81.67
3 InstructBlip 75.94
4 Qwen-VL 60.43
5 Shikra-VQA 58.61
6 LLAVA-1.5-13B 56.20
7 Shikra 56.08
8 MiniGPT-4 46.70
9 LLAVA-1.5-7B 45.20
10 Otter 38.39
(i) Pr.Att.-Food
Rank Model Score
  1 XComposer-VL 82.18
2 Blip2 82.16
3 InstructBlip 74.20
4 Qwen-VL 60.29
5 Shikra 54.03
6 LLAVA-1.5-13B 53.77
7 Shikra-VQA 52.86
8 MiniGPT-4 46.45
9 LLAVA-1.5-7B 42.07
10 Otter 39.52
(j) Pr.Att.-Plant
Rank Model Score
  1 XComposer-VL 60.37
2 LLAVA-1.5-7B 42.53
3 Blip2 42.20
4 LLAVA-1.5-13B 41.30
5 MiniGPT-4 39.18
6 Shikra 38.98
7 Qwen-VL 37.56
8 Shikra-VQA 37.47
9 InstructBlip 35.75
10 Otter 32.18
(k) Pr.Att.-Age
Rank Model Score
  1 Blip2 93.97
2 InstructBlip 80.44
3 XComposer-VL 79.56
4 LLAVA-1.5-13B 74.03
5 Shikra-VQA 63.62
6 Shikra 62.06
7 Qwen-VL 56.05
8 MiniGPT-4 51.87
9 LLAVA-1.5-7B 51.33
10 Otter 39.31
(l) Pr.Att.-Gender
Rank Model Score
  1 XComposer-VL 78.59
2 Blip2 67.94
3 InstructBlip 67.72
4 Shikra-VQA 65.00
5 Shikra 61.51
6 Qwen-VL 59.16
7 LLAVA-1.5-13B 58.33
8 LLAVA-1.5-7B 53.00
9 MiniGPT-4 45.37
10 Otter 39.91
(m) Pr.Att.-Expression
Rank Model Score
  1 XComposer-VL 62.21
2 Blip2 52.98
3 InstructBlip 48.42
4 LLAVA-1.5-13B 47.80
5 Qwen-VL 44.81
6 Shikra-VQA 42.09
7 LLAVA-1.5-7B 41.75
8 Shikra 41.07
9 MiniGPT-4 38.03
10 Otter 28.77
(n) Pr.Att.-Race
Rank Model Score
  1 Blip2 90.07
2 XComposer-VL 87.80
3 InstructBlip 82.43
4 Qwen-VL 65.20
5 LLAVA-1.5-13B 60.13
6 Shikra-VQA 55.53
7 Shikra 53.98
8 MiniGPT-4 48.02
9 Otter 46.92
10 LLAVA-1.5-7B 44.73
(o) Pr.Att.-Animal
Rank Model Score
  1 Blip2 83.55
2 XComposer-VL 81.03
3 InstructBlip 76.06
4 Qwen-VL 61.06
5 LLAVA-1.5-13B 56.84
6 Shikra-VQA 55.02
7 MiniGPT-4 49.61
8 Shikra 46.96
9 LLAVA-1.5-7B 44.26
10 Otter 38.32
(p) Pr.Att.-Object
Rank Model Score
  1 XComposer-VL 60.60
2 Blip2 52.84
3 LLAVA-1.5-13B 45.26
4 Qwen-VL 43.55
5 LLAVA-1.5-7B 42.23
6 Shikra 41.06
7 MiniGPT-4 40.32
8 Shikra-VQA 40.12
9 InstructBlip 40.03
10 Otter 33.86
(q) Pr.Att.-Style
Rank Model Score
  1 XComposer-VL 64.99
2 Shikra-VQA 55.50
3 Blip2 54.94
4 Shikra 54.16
5 LLAVA-1.5-13B 51.16
[…]

### v3

@@ Appendix A The Leaderboards
The model performance results for each subtask under clean scenario are shown in Tab. 5, Tab. 6, Tab. 7, and Tab. 8. Since the free-form question allows to assess the model’s perception abilities across multiple subtasks at the same time, the results of free-form are not taken into account.
Table 5: Evaluation results on the 5 perceptual subtasks. The top two results on each subtask are bolded and underlined, respectively. “MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
Model LLM Visual Encoder Movie Action Tv Show Profession Landmark
MC TF MC TF MC TF MC TF MC TF
MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 29.53 51.03 45.06 50.67 30.43 47.09 35.85 46.88 52.37 51.15
MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 49.76 49.20 68.99 50.16 38.24 39.47 51.85 49.02 68.67 51.69
MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 61.35 52.94 91.38 57.72 47.06 55.26 81.48 49.02 87.95 58.43
MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 72.95 50.27 95.31 62.22 73.53 52.63 79.63 62.75 87.95 50.56
BLIP2 Flan-T5-XL EVA-CLIP ViT-G 72.45 67.72 97.16 93.02 57.97 57.00 81.60 75.45 98.11 93.44
BLIP2 OPT-3B EVA-CLIP ViT-G 31.88 41.18 34.49 48.23 32.35 47.37 25.93 54.90 32.53 43.82
BLIP2 OPT-7B EVA-CLIP ViT-G 19.90 42.78 24.05 47.91 23.53 47.37 18.52 50.98 24.39 39.33
InstructBLIP Vicuna-7B EVA-CLIP ViT-G 83.57 56.68 97.88 71.38 88.24 55.26 81.48 58.82 97.59 71.91
InstructBLIP Vicuna-13B EVA-CLIP ViT-G 83.57 59.36 98.79 64.63 88.24 63.16 77.78 60.78 97.59 66.29
InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 77.93 67.24 97.60 93.98 68.12 56.04 82.08 75.00 97.16 95.08
InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 75.80 62.83 97.75 93.54 66.67 61.84 84.91 78.57 96.85 90.16
LLava-1.5 Vicuna-7B CLIP ViT-L 55.71 53.54 74.10 50.89 53.14 57.49 57.08 58.93 83.28 55.08
LLava-1.5 Vicuna-13B CLIP ViT-L 65.45 55.91 79.04 59.58 58.45 58.45 67.45 62.95 94.01 57.05
Otter LLaMA-7B CLIP ViT-L 66.01 59.53 66.62 72.44 70.05 57.00 68.87 55.36 59.31 66.56
Shikra LLaMA-7B CLIP ViT-L 68.34 61.26 78.44 78.01 60.87 57.97 82.08 68.30 88.96 70.82
Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 80.82 77.64 97.01 94.80 78.26 70.53 84.43 76.34 97.16 95.41
Xcomposer2-VL InternLM2-7B CLIP ViT-L 91.79 88.77 98.18 95.34 85.29 84.21 88.89 90.20 98.80 100.00
Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 71.08 49.61 95.96 63.52 68.12 42.51 80.66 51.34 95.90 55.41
Emu2-Chat LLaMA-33B EVA2-CLIP-E 91.30 54.55 98.79 55.14 85.29 50.00 90.74 41.18 100.00 69.66
GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 93.72 91.98 99.55 98.39 97.06 89.47 88.89 94.12 100.00 100.00
MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 93.24 89.84 98.18 92.77 94.12 89.47 90.74 88.24 100.00 98.88
Yi-VL Yi-6B-Chat OpenClip ViT-H 89.37 82.89 97.88 85.85 88.24 89.47 94.44 74.51 97.59 93.26
mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 90.34 76.47 96.82 85.37 82.35 81.58 87.04 82.35 98.80 86.52
Phi-3-Vision Phi-3 CLIP ViT-L 87.44 70.05 97.13 90.03 85.29 65.79 88.89 84.31 100.00 96.63
[…]
Gemini-1.5-Pro / / 92.72 63.64 99.24 89.55 93.55 68.42 90.74 84.31 100.00 94.38
Table 6: Evaluation results on the 5 perceptual subtasks. The top two results on each subtask are bolded and underlined, respectively. “MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
Model LLM Visual Encoder Anime Clothes Celebrity Food Plant
MC TF MC TF MC TF MC TF MC TF
MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 27.44 47.27 30.61 49.49 27.26 47.87 43.96 50.00 47.07 49.94
MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 44.25 45.14 47.76 53.57 31.06 51.05 73.66 51.21 68.35 49.94
MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 40.71 53.47 66.94 51.43 54.49 54.46 81.61 52.90 84.97 58.14
MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 50.44 47.92 65.71 47.50 57.81 53.00 86.58 56.76 89.76 59.38
BLIP2 Flan-T5-XL EVA-CLIP ViT-G 56.91 61.77 84.08 75.26 80.78 76.45 93.24 90.56 92.53 90.15
BLIP2 OPT-3B EVA-CLIP ViT-G 26.55 45.14 28.57 45.00 31.89 48.70 32.42 48.31 35.11 48.82
BLIP2 OPT-7B EVA-CLIP ViT-G 24.78 46.53 29.39 45.00 23.29 51.86 26.96 49.15 26.73 50.25
InstructBLIP Vicuna-7B EVA-CLIP ViT-G 83.19 70.14 89.80 58.21 80.40 58.67 93.66 68.12 96.28 73.29
InstructBLIP Vicuna-13B EVA-CLIP ViT-G 76.11 62.50 93.06 60.71 77.91 54.13 95.03 65.46 95.61 72.30
InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 60.77 63.65 88.27 82.72 83.11 75.75 93.96 91.36 92.98 91.36
InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 62.80 67.42 87.24 87.32 82.11 82.93 94.28 91.91 94.01 93.90
LLava-1.5 Vicuna-7B CLIP ViT-L 48.37 48.02 47.86 50.92 56.38 54.56 56.68 50.83 50.80 50.06
LLava-1.5 Vicuna-13B CLIP ViT-L 58.54 49.34 65.41 59.00 61.04 57.14 82.51 56.15 79.59 55.91
Otter LLaMA-7B CLIP ViT-L 62.80 54.99 47.24 71.37 44.15 63.76 44.44 79.46 69.29 80.56
Shikra LLaMA-7B CLIP ViT-L 47.15 58.19 76.63 59.71 63.90 62.65 84.26 67.72 88.73 66.39
Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 74.80 74.01 86.84 87.22 88.23 87.25 93.72 90.96 92.34 91.55
Xcomposer2-VL InternLM2-7B CLIP ViT-L 88.50 84.03 95.10 92.50 84.72 88.82 96.89 94.08 97.74 93.42
Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 73.37 54.61 80.71 62.17 87.17 50.03 92.05 54.24 90.99 59.21
Emu2-Chat LLaMA-33B EVA2-CLIP-E 86.73 66.67 94.69 49.64 94.02 47.16 97.76 49.15 98.54 54.66
GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 92.92 96.53 95.51 87.86 98.67 96.11 98.26 92.63 99.07 96.89
MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 82.30 89.58 91.84 92.50 88.04 80.71 96.02 93.48 98.14 95.03
Yi-VL Yi-6B-Chat OpenClip ViT-H 77.88 81.94 92.24 76.07 87.21 86.71 95.28 79.59 96.14 79.63
mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 83.19 70.14 89.80 81.79 86.05 72.61 93.79 84.18 94.95 84.47
Phi-3-Vision Phi-3 CLIP ViT-L 76.11 72.22 88.57 90.00 83.89 62.07 94.16 91.79 96.01 93.66
[…]
Gemini-1.5-Pro / / 91.82 76.39 97.55 93.57 94.85 91.57 98.01 92.75 97.87 93.54
Table 7: Evaluation results on the 5 perceptual subtasks. The top two results on each subtask are bolded and underlined, respectively. “MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
Model LLM Visual Encoder Age Gender Expression Race Animal
MC TF MC TF MC TF MC TF MC TF
MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 30.02 48.28 52.00 46.75 42.29 49.82 25.10 48.42 45.54 49.39
MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 29.32 50.06 58.30 48.60 59.51 53.58 33.08 51.51 60.39 52.37
MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 49.48 49.74 84.34 56.14 57.26 57.74 28.35 52.96 89.84 54.49
MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 41.62 52.84 86.13 61.60 65.69 63.85 51.34 53.56 92.53 60.20
BLIP2 Flan-T5-XL EVA-CLIP ViT-G 62.72 58.71 99.32 94.12 88.41 75.30 75.76 71.87 96.62 95.59
BLIP2 OPT-3B EVA-CLIP ViT-G 24.61 43.28 61.41 48.60 34.53 53.84 31.55 51.51 36.92 48.78
BLIP2 OPT-7B EVA-CLIP ViT-G 29.84 42.76 53.13 47.81 30.93 53.78 26.55 49.82 28.40 49.43
InstructBLIP Vicuna-7B EVA-CLIP ViT-G 70.16 57.36 99.66 74.00 83.13 67.23 84.16 67.55 97.46 69.17
InstructBLIP Vicuna-13B EVA-CLIP ViT-G 67.28 63.44 99.33 65.49 86.28 74.38 81.23 56.45 97.01 71.45
InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 64.06 58.48 99.58 86.91 91.11 78.06 78.28 73.98 97.08 95.52
InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 67.84 81.16 99.58 98.50 82.56 80.11 82.46 83.55 97.32 95.71
LLava-1.5 Vicuna-7B CLIP ViT-L 37.74 55.40 53.62 49.89 63.49 51.34 42.55 51.34 49.45 52.18
LLava-1.5 Vicuna-13B CLIP ViT-L 49.29 59.78 98.60 83.59 70.12 58.87 70.81 63.45 86.12 58.74
Otter LLaMA-7B CLIP ViT-L 36.78 50.92 78.99 77.84 72.50 61.02 42.38 58.62 81.57 83.57
Shikra LLaMA-7B CLIP ViT-L 65.07 54.96 98.14 73.29 90.08 70.52 75.20 54.19 90.20 70.32
Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 67.09 79.02 99.53 97.57 90.25 81.34 79.47 76.26 97.43 96.32
Xcomposer2-VL InternLM2-7B CLIP ViT-L 86.65 91.73 99.11 99.03 96.74 94.41 88.38 89.02 98.06 89.56
Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 53.19 48.71 97.50 59.18 85.55 64.04 74.43 55.95 95.45 63.89
Emu2-Chat LLaMA-33B EVA2-CLIP-E 84.95 49.35 99.89 50.91 92.58 53.06 91.95 51.39 98.21 53.51
GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 86.13 93.80 100.00 100.00 98.09 94.28 94.51 94.69 99.40 99.02
MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 77.23 68.86 99.55 99.76 95.05 92.07 81.61 88.66 98.65 94.62
[…]